# Vital Insight: Assisting Experts’ Context-Driven Sensemaking of Multi-modal Personal Tracking Data Using Visualization and Human-In-The-Loop LLM

### v3

@@ 1. Introduction
The ubiquitous presence of sensor-rich smartphones and wearables has prompted researchers to use data from these devices to track living activities for various outcomes including health tracking, context-aware applications, and digital health interventions (friedewald2011ubiquitous, ; orwat2008towards, ; mishra2018investigating, ; ho2005using, ; kunzler2019exploring, ; boateng2019experience, ; farber2011aging, ; li2023privacy, ; kononova2019use, ; vargemidis2020wearable, ; seifert2017use, ). However, longstanding socio-technical gaps persist in collecting data and generating valuable insights, due to challenges such as limited information, noisy data, ambiguous contexts, and variations in individual behavior (ackerman2000intellectual, ; chen2002context, ; vcas2011ubiquitous, ; edwards2001home, ; bourguet2011uncertainty, ; kumar2009challenges, ). Most efforts in this area focus on developing algorithms to model windows of raw sensor data into labels of activity, behavior, or outcomes, such as predicting physical activity, stress, and depression (10466375, ; amin2023investigating, ; mishra2020continuous, ; cornacchia2016survey, ; doherty2014tracking, ; wang2018tracking, ). There is however, a lack of exploration in generating high-level insights about these periodic behaviors such as “is it a normal day” or a summary of someone’s day that is often more valuable for various stakeholders (mynatt2001digital, ; adler2024beyond, ). In real-world deployment, researchers often have to find alternate ways to derive insights that are needed for practical applications.
@@ 2.2. Tools to assist the sensemaking of personal tracking data
Some past research has acknowledged the challenges of making sense of personal tracking data and has developed various tools to assist with different tasks. Among those tools aimed at experts, many focus on specific tasks, such as activity annotation on accelerometer data (ponnada2021signaligner, ; hagen2023annots, ), multi-modal sensor data (barz2016multimodal, ; martindale2018smart, ; diete2017smart, ), video analysis (10.1145/2968219.2971461, ; diete2017recognizing, ; sztyler2016body, ), and spatial-temporal events (rawassizadeh2019indexing, ). The majority of these tools are built around visualization, where the system serves as an interactive interface to better display data and assist in the annotation process. Beyond activity annotation, some tools take a broader approach to understanding multi-modal data. For instance, Lifestream integrates different data modalities and features into one dashboard to facilitate the exploration and evaluation of personal data streams with users (hsieh2013lifestreams, ). Chung et al. used the EventFlow tool, which aggregates workflow information to highlight patterns and variabilities, helping different stakeholders understand the daily routines of older adults (chung2017examining, ). Some works also explored other modalities, such as conversational interfaces in helping the understanding of data (kim2021data, ; shen2022towards, ; hearst2019would, ; aurisano2015show, ).
@@ 3.1. Method
We conducted an IRB-approved formative interviews with 12 experts in multi-modal sensing data with a \$15 compensation. The inclusion criteria for experts are experienced researchers in ubiquitous computing who have utilized passive sensing data for health-related applications. In addition to self-selection, we also added objective inclusion criteria, requiring publication in IMWUT, CHI, JMIR, and other relevant conferences or journals. Through these hour-long semi-structured online interviews, we wanted to better understand the experts’ unique research processes while inferring complex real-world personal tracking data. After obtaining verbal consent, we recorded the audio and transcripts of the interviews. Two members of the research team carefully reviewed the transcripts and notes, and independently coded specific segments of the data, employing an inductive open coding approach (corbin1990grounded, ). The coders then exchanged codebooks and discussed to harmonize and merge the codes by consensus, and conducted a thematic analysis under the principles of Grounded Theory (braun2006using, ; glaser1968discovery, ). We recruited a total of 12 experts who had various experiences in working with personal tracking data (Table 1), reached data saturation and stopped further recruitment (Guest2006HowMI, ; Ando2014AchievingSI, ).
@@ 3.2.1. Challenging sensemaking scenarios in current practices
Another way to obtain interpretation of data is to source directly from participants. However, discussions with experts reveal another significant challenge: understanding and utilizing self-report data. Self-report data is often limited, and experts must be cautious in treating self-report data as absolute truth, especially when dealing with older adults or those with cognitive impairments (P4). Even when experts manage to obtain some level of ground truth, they still face the challenge of integrating this information with other data modalities, mirroring the process of contextualizing raw sensor data. This gap makes it difficult for experts to generate and validate their interpretations.
@@ 3.2.2. Needs and Opportunities in assisting experts
The first approach of examine the sensing data to identify any emerging patterns or insights allows experts to uncover unexpected trends and behaviors that might not have initially considered. For instance, P8 remarked, “it wasn’t until we got into the project that we realized maybe there’s other aspects we need to be looking at.” P6 likened the process to a “fishing expedition,” explaining that “more people are always coming up with new and interesting ways of extrapolating info out of this data.”(P6). Starting with the data enables experts to discover novel, previously undocumented insights and generate ideas that could lead to a deeper understanding of the data or behaviors of the observed participants.
@@
Now, recognizing the iterative nature of the sensemaking process and the two distinct paths it can take, we must consider how we can support this process effectively. We have identified two primary needs based on these paths: direct representation of data and indirect inference. Experts require a direct representation of sensing data to manage the vast volumes of information, as sensor data often does not come in a human-readable format. Throughout our interviews, experts commonly relied on visualizations for this purpose. However, most visualizations lack systematic design, are overly generic, and are often created ad-hoc for specific tasks. Experts mentioned techniques like stacking modalities or creating correlation graphs, but these efforts are sporadic and have not been standardized even within the same research group. Beyond direct data representation, participants also expressed a need for assistance with automatic inference generation as potential guidance for their sensemaking. Even with effective visualizations, the transformation of data into actionable insights is still predominantly a manual process that the experts conduct, which is overwhelming when dealing with so much data (P9). In longitudinal studies that span months or even years, it becomes impractical for experts to examine every piece of data to get the desired granularity required for some applications like labels for machine learning (P12/13). As a result, assistance in quickly navigating to the desired information seemed crucial. Our participants noted a significant lack of tools for this purpose, as the sensemaking process is complex and often beyond the capabilities of current algorithms and tools. For instance, some experts who used machine learning in their studies mentioned that they only viewed benchmarks as “results to be compared” against ground truth data, rather than as useful assistance for them. However, they expressed interest in getting some extracted data insights that go beyond simple labels. Based on these insights, we designed and developed our initial prototype.
@@ 4.1. Data types and system infrastructure
Figure 1. Eight components in VI-1 powered by visualization and LLM: time selection, main visualization, user profile, user check-in (user conversation with chatbot), LLM daily summary, LLM question, LLM anomaly detection, LLM hourly summary
[…]
Screenshot of the Vital Insight interface showing the User Profile on the left with details like gender, education, race, marital status, occupation, living arrangement, addresses, recreation activities. On the right, two Anomaly panels display heart rate and respiration issues, with timestamps and explanations for spikes in heart rate and increased respiration related to Rapid Eye Movement sleep.
@@ 4.2. Interface design
Our initial prototype has two major parts: (1) visualizations that provide direct representation of sensor data, and (2) varying granularities of summaries and inferences generated by LLM that provides indirect inference. Our system interface has five panels with eight components: time selection, main visualization, user profile, daily summary (user check-in, LLM-generated daily summary and questions), LLM-anomaly, and LLM-hourly summary (Fig. 1). To generate the interface, our system uses data from a real study with participants, which we discuss in greater detail in Sec. 4.3.
@@
The Main Visualization panel serves as the central component to represent the sensing data (Fig.  1 left). The visualization component features five categories of information gathered from the phone, smartwatch, and chatbot: location, health, phone usage, activity, and events. We used a small multiple-style time series plot arrangement to display each category of information along a unified timeline. The design also aligns with current practices, where experts manually stack different modalities of data, as described in the interview (Section 3.2.2). The Location in the background shows location labels, e.g., home and church. We generate this data by cross-referencing the raw GPS coordinates from the phone app with participants’ self-reported frequently visited locations and addresses. The system marks each location with a unique background color across all views with a label at the top, aiming to provide contextual information when experts examine the data. In Fig. 1, the plotted location labels are home (light green) and church (light orange). The Health view presents heart rate and respiration data from the smartwatch using scatter plots in red and blue. Two black trend-lines overlayed on the graph represent the participant’s average heart rate and respiration over a one-month period. Exclamation marks (!) next to the scatter plot indicate detected anomalies. We present more information about the anomaly detection module in section 4.3. The Phone Usage view includes Wi-Fi (connected/disconnected, purple line chart), phone battery levels (0-100, green line chart), and phone interaction (unlock/lock events, solid/dashed grey rectangle overlaid). The Activity view reports step counts and basic activity recognition. A bar chart presents the step counts, with the height of the bar representing the number of steps taken during the time indicated by the width. The activity view presents basic activities from iPhone’s standard activity recognition API, using different colored rectangles (‘stationary’ in gray, ’walking’ in green, ’automotive’ in purple), indicating the activities and their respective time spans. Finally, the Event view includes significant events throughout the day, including phone calls (green) and interactions with the Alexa chatbot (orange), and their respective durations.
@@
Daily Summary panel, provides a high-level overview of the participant’s day with three tabs: LLM Summary, User Check-in, and Questions. The LLM Summary tab presents an LLM-generated summary of the entire day in paragraphs and bullet points for easy reading, thus providing experts the ability to quickly understand a person’s day. We selected LLM to generate the summary due to its extensive capability to customize text-format summaries and explanations, offering different levels of detail and explainability for experts. The User Check-in tab displays the real conversation between the chatbot and the participant (orange ‘Chatbot’ event in visualization). The Questions tab presents a list of questions and missing information that the system identifies as potentially helpful for data interpretation. This information can serve as an inspiration for experts to reflect on potential actions, particularly when engaging with different stakeholders, such as speaking directly with participants (Scenario 4). We present detailed rationale for the LLM-generated content in Sec. 4.3.
@@
The Anomaly panel provides insights about anomalies and outliers, typically defined as “an observation that deviates significantly from other observations, raising suspicions that it was generated by a different mechanism,” and is crucial in analyzing time-series data such as heart rate (hawkins1980identification, ). In previous interviews, experts noted that anomalies are worth paying attention to (Scenario 3). For each anomaly, the panel displays the specific time range, description, and a possible reason, all generated by the LLM. Users can select the specific anomaly time to view the detailed information. The Hourly Summary panel provides LLM-generated interpretations for each hour of data. Users can select a specific time (e.g. 4-5 AM) to view the corresponding interpretation. Similar to the daily summary, this panel offers an objective summary of the sensor data with minimal inference, a more creative insights derived from the data, and questions that the LLM identifies as useful for interpreting that hour’s data. This panel offers different levels of inferences on potential events in detail (Scenario 2).
@@
Additionally, we implemented several interactive functions to improve the system’s usability. To enhance the visual alignment of different modalities, we added a vertical tooltip that appears across all modalities when the user hovers over the visualization and displays the time and value. This helps experts easily compare and integrate different sensing modalities for interpretation. To avoid label overlap on a sequence of narrow rectangular boxes of phone unlock/lock and activity, we implemented a hover effect that reveals labels (e.g., “walking”) only when a user hovers over it.
@@ 4.3. LLM Augmentation
There are two primary forms of data input: Sensing Data and Contextual Data. For Sensing Data, we process the data into semantic description before sending it to the model. Discrete data, such as WiFi and phone battery levels, are directly converted into sentences with values and timestamps (e.g. “The battery level of the person’s phone is [BATTERY] at [TIME].”). For time-series data like heart rate, encoding each data point individually would overwhelm our system. Instead, since these data are typically sampled at a fixed rate, we group them and send them in an array format like “The person’s respiration from [TIME] to [TIME] (10-second interval) is [VALUE1, VALUE2, …].”, and sort this narrative in chronological order. Another data input is Contextual data which includes User Profile - demographic and routine information, and User Check-in - which we process conversation in a time(From [TIME] to [TIME]) and utterance ([ROLE]:[UTTERANCE]) format. We do not specifically handle missing or noisy data and transfer each data point as-is.
Figure 2. Overview of the structure of LLM augmentation for VI-1 and VI-2.
[…]
Diagram showing the structure of the LLM system with two main sections: Input and Output in Vital Insight. The Input section includes ’Sensor Data’ labeled as Semantic Sensing Data and ’Contextual Information’ with User Profile, User Check-in, and Daily Feature. Arrows connect these inputs to the Output section, which features ’Hourly Data’ with tabs for Summary, Inference, and Question. Outputs from Hourly Data lead to ’Daily Summary’ for LLM Summary, ’Question’ for high-level questions, and ’Anomaly’ for detection and explanation.
@@
The prototype then processes the low-level summary into LLM High-Level sensemaking. Recognizing the daily patterns in activity, it aggregates the outputs from 24 hours of the hourly summary and other contextual information, to generate a similar output containing a summary, inference, and questions, but on a daily basis. The prompt includes the Goal (“You are tasked with summarizing a user’s daily activities based on the provided information…”), User Profile, User Check-in, Historical Summary from GPT (24-hour output from the hourly function), and the JSON Formatter. The system generates and formats the results in a human-readable way, both as a paragraph and in bullet points for easier digestion. It also summarizes information like questions at a higher daily level over a 24-hour period. Before sending the summary to the dashboard, we perform several programmatic checks and rerun the process if the output does not meet the standard (e.g., output tokens¡10).
Other than summary from the hourly data, LLM anomaly detector identifies potential anomalies in different modalities directly using the semantic descriptions of each sensing modality. While we iteratively adjusted these prompts to optimize outputs, we provided all LLM-generated results as-is and unaltered during user testing. If the inputs exceed token limits, we split them into chunks before sending them to the model; however, after this multi-level procedure, it’s very unlikely that the input tokens will exceed the context window of the model (128k).

### v4

@@ Abstract.
AI-assisted Sensemaking, Personal Tracking, Large Language Models, Thematic Analysis
copyright: ccjournal: IMWUTjournalyear: 2025journalvolume: 9journalnumber: 3article: 101publicationmonth: 9doi: 10.1145/3749508ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Artificial intelligence
@@ 1. INTRODUCTION
The ubiquitous presence of sensor-rich smartphones and wearables has prompted researchers to use data from these devices to track living activities for various outcomes including health tracking, context-aware applications, and digital health interventions (friedewald2011ubiquitous, ; orwat2008towards, ; mishra2018investigating, ; ho2005using, ; kunzler2019exploring, ; boateng2019experience, ; farber2011aging, ; li2023privacy, ; kononova2019use, ; vargemidis2020wearable, ; seifert2017use, ; 10.1145/3706598.3714086, ). However, longstanding socio-technical gaps persist in collecting data and generating valuable insights, due to challenges such as limited information, noisy data, ambiguous contexts, and variations in individual behavior (ackerman2000intellectual, ; chen2002context, ; vcas2011ubiquitous, ; edwards2001home, ; bourguet2011uncertainty, ; kumar2009challenges, ). Most efforts in this area focus on developing algorithms to model windows of raw sensor data into labels of activity, behavior, or outcomes, such as predicting physical activity, stress, and depression (10466375, ; amin2023investigating, ; mishra2020continuous, ; cornacchia2016survey, ; doherty2014tracking, ; wang2018tracking, ). There is however, a lack of exploration in generating high-level insights about these periodic behaviors such as “is it a normal day” or a summary of someone’s day that is often more valuable for various stakeholders (mynatt2001digital, ; adler2024beyond, ). In real-world deployment, researchers often have to find alternate ways to derive insights that are needed for practical applications.
@@ 2.2. Tools to Assist the Sensemaking of Personal Tracking Data
Some past research has acknowledged the challenges of making sense of personal tracking data and has developed various tools to assist with different tasks. Among those tools aimed at experts, many focus on specific tasks, such as activity annotation on accelerometer data (ponnada2021signaligner, ; hagen2023annots, ), multi-modal sensor data (barz2016multimodal, ; martindale2018smart, ; diete2017smart, ), video analysis (10.1145/2968219.2971461, ; diete2017recognizing, ; sztyler2016body, ), and spatial-temporal events (rawassizadeh2019indexing, ). The majority of these tools are built around visualization, where the system serves as an interactive interface to better display data and assist in the annotation process. Beyond activity annotation, some tools take a broader approach to understanding multi-modal data. For instance, Lifestream integrates different data modalities and features into one dashboard to facilitate the exploration and evaluation of personal data streams with users (hsieh2013lifestreams, ). Chung et al. used the EventFlow tool, which aggregates workflow information to highlight patterns and variabilities, helping different stakeholders understand the daily routines of older adults (chung2017examining, ). Some works also explored other modalities, such as conversational interfaces in helping the understanding of data (kim2021data, ; shen2022towards, ; hearst2019would, ; aurisano2015show, ; 10.1145/3706598.3714272, ).
@@ 3.1. Method
We conducted IRB-approved formative interviews with 12 experts in multi-modal sensing data with a \$15 compensation. The inclusion criteria for experts are experienced researchers in ubiquitous computing who have utilized passive sensing data for health-related applications. In addition to self-selection, we also added objective inclusion criteria, requiring publication in IMWUT, CHI, JMIR, and other relevant conferences or journals. Through these hour-long semi-structured online interviews, we wanted to better understand the experts’ unique research processes while inferring complex real-world personal tracking data. After obtaining verbal consent, we recorded the audio and transcripts of the interviews. Two members of the research team carefully reviewed the transcripts and notes, and independently coded specific segments of the data, employing an inductive open coding approach (corbin1990grounded, ). The coders then exchanged codebooks and discussed to harmonize and merge the codes by consensus, and conducted a thematic analysis under the principles of Grounded Theory (braun2006using, ; glaser1968discovery, ). We recruited a total of 12 experts who had various experiences in working with personal tracking data (Table 1), reached data saturation and stopped further recruitment (Guest2006HowMI, ; Ando2014AchievingSI, ).
@@ 3.2.1. Challenging Sensemaking Scenarios in Current Practices
Another way to obtain an interpretation of data is to source directly from participants. However, discussions with experts reveal another significant challenge: understanding and utilizing self-report data. Self-report data is often limited, and experts must be cautious in treating self-report data as absolute truth, especially when dealing with older adults or those with cognitive impairments (P4). Even when experts manage to obtain some level of ground truth, they still face the challenge of integrating this information with other data modalities, mirroring the process of contextualizing raw sensor data. This gap makes it difficult for experts to generate and validate their interpretations.
@@ 3.2.2. Needs and Opportunities in Assisting Experts
The first approach to examine the sensing data to identify any emerging patterns or insights allows experts to uncover unexpected trends and behaviors that might not have been initially considered. For instance, P8 remarked, “it wasn’t until we got into the project that we realized maybe there are other aspects we need to be looking at.” P6 likened the process to a “fishing expedition,”, explaining that “more people are always coming up with new and interesting ways of extrapolating info out of this data.”(P6). Starting with the data enables experts to discover novel, previously undocumented insights and generate ideas that could lead to a deeper understanding of the data or behaviors of the observed participants.
@@
Now, recognizing the iterative nature of the sensemaking process and the two distinct paths it can take, we must consider how we can support this process effectively. We have identified two primary needs based on these paths: direct representation of data and indirect inference. Experts require a direct representation of sensing data to manage the vast volumes of information, as sensor data often does not come in a human-readable format. Throughout our interviews, experts commonly relied on visualizations for this purpose. However, most visualizations lack systematic design, are overly generic, and are often created ad hoc for specific tasks. Experts mentioned techniques like stacking modalities or creating correlation graphs, but these efforts are sporadic and have not been standardized even within the same research group. Beyond direct data representation, participants also expressed a need for assistance with automatic inference generation as potential guidance for their sensemaking. Even with effective visualizations, the transformation of data into actionable insights is still predominantly a manual process that the experts conduct, which is overwhelming when dealing with so much data (P9). In longitudinal studies that span months or even years, it becomes impractical for experts to examine every piece of data to get the desired granularity required for some applications like labels for machine learning (P12/13). As a result, assistance in quickly navigating to the desired information seemed crucial. Our participants noted a significant lack of tools for this purpose, as the sensemaking process is complex and often beyond the capabilities of current algorithms and tools. For instance, some experts who used machine learning in their studies mentioned that they only viewed benchmarks as “results to be compared” against ground truth data, rather than as useful assistance for them. However, they expressed interest in getting some extracted data insights that go beyond simple labels. Based on these insights, we designed and developed our initial prototype.
@@ 4.1. Data Types and System Infrastructure
Figure 1. Eight components in VI-1 powered by visualization and LLM: time selection, main visualization, user profile, user check-in (user conversation with chatbot), LLM daily summary, LLM question, LLM anomaly detection, LLM hourly summary:
@@ 4.2. Interface Design
Our initial prototype has two major parts: (1) visualizations that provide direct representation of sensor data, and (2) varying granularities of summaries and inferences generated by LLM that provide indirect inference. Our system interface has five panels with eight components: time selection, main visualization, user profile, daily summary (user check-in, LLM-generated daily summary and questions), LLM-anomaly, and LLM-hourly summary (Fig. 1). To generate the interface, our system uses data from a real study with participants, which we discuss in greater detail in Sec. 4.3.
@@
The Main Visualization panel serves as the central component to represent the sensing data (Fig.  1 left). The visualization component features five categories of information gathered from the phone, smartwatch, and chatbot: location, health, phone usage, activity, and events. We used a small multiple-style time series plot arrangement to display each category of information along a unified timeline. The design also aligns with current practices, where experts manually stack different modalities of data, as described in the interview (Section 3.2.2). The Location in the background shows location labels, e.g., home and church. We generate this data by cross-referencing the raw GPS coordinates from the phone app with participants’ self-reported frequently visited locations and addresses. The system marks each location with a unique background color across all views with a label at the top, aiming to provide contextual information when experts examine the data. In Fig. 1, the plotted location labels are home (light green) and church (light orange). The Health view presents heart rate and respiration data from the smartwatch using scatter plots in red and blue. Two black trend lines overlayed on the graph represent the participant’s average heart rate and respiration over a one-month period. Exclamation marks (!) next to the scatter plot indicate detected anomalies. We present more information about the anomaly detection module in section 4.3. The Phone Usage view includes Wi-Fi (connected/disconnected, purple line chart), phone battery levels (0-100, green line chart), and phone interaction (unlock/lock events, solid/dashed grey rectangle overlaid). The Activity view reports step counts and basic activity recognition. A bar chart presents the step counts, with the height of the bar representing the number of steps taken during the time indicated by the width. The activity view presents basic activities from iPhone’s standard activity recognition API, using different colored rectangles (‘stationary’ in gray, ’walking’ in green, ’automotive’ in purple), indicating the activities and their respective time spans. Finally, the Event view includes significant events throughout the day, including phone calls (green) and interactions with the Alexa chatbot (orange), and their respective durations.
@@
Daily Summary panel, provides a high-level overview of the participant’s day with three tabs: LLM Summary, User Check-in, and Questions. The LLM Summary tab presents an LLM-generated summary of the entire day in paragraphs and bullet points for easy reading, thus providing experts the ability to quickly understand a person’s day. We selected LLM to generate the summary due to its extensive capability to customize text-format summaries and explanations, offering different levels of detail and explainability for experts. The User Check-in tab displays the real conversation between the chatbot and the participant (orange ‘Chatbot’ event in visualization). The Questions tab presents a list of questions and missing information that the system identifies as potentially helpful for data interpretation. This information can serve as an inspiration for experts to reflect on potential actions, particularly when engaging with different stakeholders, such as speaking directly with participants (Scenario 4). We present detailed rationales for the LLM-generated content in Sec. 4.3.
@@
The Anomaly panel provides insights about anomalies and outliers, typically defined as “an observation that deviates significantly from other observations, raising suspicions that it was generated by a different mechanism,” and is crucial in analyzing time-series data such as heart rate (hawkins1980identification, ). In previous interviews, experts noted that anomalies are worth paying attention to (Scenario 3). For each anomaly, the panel displays the specific time range, description, and a possible reason, all generated by the LLM. Users can select the specific anomaly time to view the detailed information. The Hourly Summary panel provides LLM-generated interpretations for each hour of data. Users can select a specific time (e.g. 4-5 AM) to view the corresponding interpretation. Similar to the daily summary, this panel offers an objective summary of the sensor data with minimal inference, more creative insights derived from the data, and questions that the LLM identifies as useful for interpreting that hour’s data. This panel offers different levels of inferences on potential events in detail (Scenario 2).
@@
Additionally, we implemented several interactive functions to improve the system’s usability. To enhance the visual alignment of different modalities, we added a vertical tooltip that appears across all modalities when the user hovers over the visualization and displays the time and value. This helps experts easily compare and integrate different sensing modalities for interpretation. To avoid label overlap on a sequence of narrow rectangular boxes of phone unlock/lock and activity, we implemented a hover effect that reveals labels (e.g., “walking”) only when a user hovers over them.
@@ 4.3. LLM Augmentation
There are two primary forms of data input: Sensing Data and Contextual Data. For Sensing Data, we process the data into semantic description before sending it to the model. Discrete data, such as WiFi and phone battery levels, are directly converted into sentences with values and timestamps (e.g. “The battery level of the person’s phone is [BATTERY] at [TIME].”). For time-series data like heart rate, encoding each data point individually would overwhelm our system. Instead, since these data are typically sampled at a fixed rate, we group them and send them in an array format like “The person’s respiration from [TIME] to [TIME] (10-second interval) is [VALUE1, VALUE2, …].”, and sort this narrative in chronological order. Another data input is Contextual data, which includes User Profile - demographic and routine information, and User Check-in - where we process conversation in a time(From [TIME] to [TIME]) and utterance ([ROLE]:[UTTERANCE]) format. We do not specifically handle missing or noisy data and transfer each data point as-is.
Figure 2. Overview of the structure of LLM augmentation for VI-1 and VI-2.:
@@
The prototype then processes the low-level summary into LLM High-Level sensemaking. Recognizing the daily patterns in activity, it aggregates the outputs from 24 hours of the hourly summary and other contextual information to generate a similar output containing a summary, inference, and questions, but on a daily basis. The prompt includes the Goal (“You are tasked with summarizing a user’s daily activities based on the provided information…”), User Profile, User Check-in, Historical Summary from GPT (24-hour output from the hourly function), and the JSON Formatter. The system generates and formats the results in a human-readable way, both as a paragraph and in bullet points for easier digestion. It also summarizes information like questions at a higher daily level over a 24-hour period. Before sending the summary to the dashboard, we perform several programmatic checks and rerun the process if the output does not meet the standard (e.g., output tokens¡10).
Other than summaries from the hourly data, LLM anomaly detector identifies potential anomalies in different modalities directly using the semantic descriptions of each sensing modality. While we iteratively adjusted these prompts to optimize outputs, we provided all LLM-generated results as-is and unaltered during user testing. If the inputs exceed token limits, we split them into chunks before sending them to the model; however, after this multi-level procedure, it’s very unlikely that the input tokens will exceed the context window of the model (128k).